# The Law of Multi-Model Collaboration: Scaling Limits of Model Ensembling for Large Language Models

### v1

@@ Abstract
Recent advances in large language models (LLMs) have been largely driven by scaling laws for individual models, which predict performance improvements as model parameters and data volume increase. However, the capabilities of any single LLM are inherently bounded. One solution originates from intricate interactions among multiple LLMs, rendering their collective performance surpass that of any constituent model. Despite the rapid proliferation of multi-model integration techniques such as model routing and post-hoc ensembling, a unifying theoretical framework of performance scaling for multi-model collaboration remains absent. In this work, we propose the Law of Multi-model Collaboration, a scaling law that predicts the performance limits of LLM ensembles based on their aggregated parameter budget. To quantify the intrinsic upper bound of multi-model collaboration, we adopt a method-agnostic formulation and assume an idealized integration oracle where the total cross-entropy loss of each sample is determined by the minimum loss of any model in the model pool. Experimental results reveal that multi-model systems follow a power-law scaling with respect to the total parameter count, exhibiting a more significant improvement trend and a lower theoretical loss floor compared to single model scaling. Moreover, ensembles of heterogeneous model families achieve better performance scaling than those formed within a single model family, indicating that model diversity is a primary driver of collaboration gains. These findings suggest that model collaboration represents a critical axis for extending the intelligence frontier of LLMs.
@@ 1 Introduction
Motivated by these constraints, various multi-model integration approaches have been proposed (Chen et al., 2025b)—including dynamic routing (Qian et al., 2025), ensemble voting (Zhao et al., 2025), tool orchestration (Zhang et al., 2025), and agent-based collaboration (Tran et al., 2025; Chen et al., 2025a; Li et al., 2024; Fan et al., 2025)—which demonstrate superior performance and efficiency over individual models across diverse application scenarios. But these successes remain empirical and method-specific, justified through heuristics or task evaluations rather than principled understanding of why multi-model collaboration works. The result is a serious mismatch between theory and practice: although the scaling law explains the growth behavior of individual models, there is no similar framework for multi-model collaboration. Does the multi-model collaborative system scale with the total number of parameters in the model pool? Can the performance upper limit of a multi-model collaborative system be predicted based on the total number of parameters in the model pool? Besides the total number of parameters in the model pool, what other factors affect the performance of multi-model collaborative systems? Without theoretical grounding, we cannot compare multi-model systems to large single models rigorously or reason about their ultimate limits. This gap motivates a general law for multi-model collaboration, parallel to existing single-model scaling laws (Kaplan et al., 2020; Hoffmann et al., 2022; Kim et al., 2025).
@@
We propose the Law of Multi-model Collaboration, which characterizes the performance limit of multi-model collaborative systems as a function of the total number of parameters of all models involved. Rather than focusing on specific ensemble mechanisms, we adopt a method-agnostic view treating multi-model systems as collective entities. This study operates within the AI Flow framework (An et al., 2025; Shao and Li, 2025), which advocates distributed intelligence integrating artificial intelligence with communication systems through collaborative interactions across heterogeneous models deployed at device, edge, and cloud levels (Yuan et al., 2025). Understanding fundamental performance limits of multi-model collaboration becomes prerequisite for principled system design in this paradigm. Our formulation centers on a model pool of LLMs with total parameter count summed across the pool. For a given input distribution, we characterize the minimal achievable loss through model collaboration, defining a theoretical lower bound analogous to single-model scaling laws. To isolate intrinsic collaboration limits, we consider an idealized oracle integration where per-token cross-entropy loss equals the minimum achieved by any pool model. While unrealizable in practice, this oracle parallels optimality assumptions in prior scaling studies that abstract away algorithmic inefficiencies to reveal fundamental trends. We construct empirical performance frontiers by enumerating model combinations from a fixed pool, evaluating aggregated parameter budgets against achieved loss, and extracting Pareto-optimal points representing the theoretical performance envelope.
@@
Our analysis yields two findings. The multi-model system exhibits a clear power-law scaling between loss and total parameter count, similar to the single model scaling law, but with a steeper downward trend and significantly lower asymptotic loss lower bound, indicating that it utilizes parameters more effectively than the single model scaling law. Additionally, scaling efficiency depends strongly on model diversity: ensembles from different families (Yang et al., 2025; Grattafiori et al., 2024; Team et al., 2025) consistently outperform single-family ensembles under equivalent parameter budgets, suggesting architectural and training heterogeneity drives collaborative gains beyond parameter count. These findings imply model diversity constitutes a fundamental scaling dimension complementary to parameters, data, and compute. The Law of Multi-model Collaboration thus reframes ensemble methods from ad-hoc techniques into a systematic, theoretically grounded pathway for extending artificial intelligence frontiers. This work makes four contributions:
@@ 2.2 Text-level Loss under Heterogeneous Tokenization
We emphasize that we deliberately use the summed loss rather than the mean loss, since averaging would implicitly normalize by tokenizer-dependent sequence length and obscure genuine differences in modeling efficiency. The summed formulation ensures that each model is evaluated on its own induced token sequence while preserving comparability at the level of complete texts. The expected loss of model $M_{i}$ over the distribution $\mathcal{D}$ is then given by
$\mathcal{L}(M_{i})\;=\;\mathbb{E}_{x\sim\mathcal{D}}\left[\mathcal{L}(M_{i},x)\right].$ (2)
@@ 2.3 Oracle Multi-model Collaboration
That is, for each individual text, the oracle selects the single model within the pool that yields the lowest text-level loss under its own tokenizer. The expected oracle loss over the distribution $\mathcal{D}$ is
$\mathcal{L}_{\mathrm{oracle}}(\mathcal{S})\;=\;\mathbb{E}_{x\sim\mathcal{D}}\left[\mathcal{L}_{\mathrm{oracle}}(\mathcal{S},x)\right].$ (4)
@@ 2.6 Law of Multi-model Collaboration
This law generalizes classical single-model scaling laws to the multi-model setting. Crucially, both the exponent $\alpha$ and the loss floor $\mathcal{L}_{\infty}$ depend on the composition of the model pool, with heterogeneous model pools exhibiting systematically larger exponents and lower asymptotic loss than homogeneous ones.
@@ 3.2 Evaluation Dataset and Loss Measurement
All models are evaluated on a fixed held-out text dataset sampled from the distribution $\mathcal{D}$. For each model $M_{i}\in\mathcal{M}$ and each input text $x$, we compute the text-level loss $\mathcal{L}(M_{i},x)$ as defined in Equation (1), namely the sum of token-level cross-entropy losses under the model’s own tokenizer. The expected loss $\mathcal{L}(M_{i})$ is estimated by averaging $\mathcal{L}(M_{i},x)$ over the evaluation dataset. This procedure yields a single-model performance baseline that is directly comparable across models despite heterogeneous tokenization schemes.
@@ 3.3 Enumeration of Multi-model Combinations
To explore the space of multi-model collaboration, we enumerate all unordered pairwise combinations of models from the pool $\mathcal{M}$. For each pair $\mathcal{S}=\{M_{i},M_{j}\}$, we compute:
  • the aggregated parameter budget $P_{\mathrm{total}}(\mathcal{S})=P_{i}+P_{j}$, and
  • the oracle ensemble loss $\mathcal{L}_{\mathrm{oracle}}(\mathcal{S})$ as defined in Equation (3).
Each model pair is further categorized into one of two groups:
@@
    Homogeneous combinations, where both models belong to the same model family;
@@
    Heterogeneous combinations, where the models originate from different model families.
This partition enables a controlled comparison of scaling efficiency with and without architectural and training diversity.
@@ 3.4 Pareto-optimal Frontier Extraction
Due to the large number of models and the wide variance in their capabilities, the raw set of evaluated points contains many dominated configurations. To focus on the intrinsic performance limits, we extract Pareto-optimal frontiers separately for:
  • single-model configurations ($\mathcal{S}=\{M_{i}\}$),
  • homogeneous multi-model combinations, and
  • heterogeneous multi-model combinations.
A configuration $\mathcal{S}$ is considered Pareto-optimal if no other configuration achieves both a lower oracle loss and a lower aggregated parameter budget. The resulting Pareto frontiers represent the empirical lower envelopes of achievable loss under the oracle collaboration assumption.
Figure 1: Comparison between scaling performance under two settings: single model and multi-model ensemble. Solid curves denote nonlinear least-squares fits of the scaling formula, while points correspond to Pareto-optimal lower envelopes. Model ensembling requires fewer parameters to achieve a target loss, and reduces the lower bound of the test loss significantly.
@@ 4.1 Scaling Behavior of Single and Multi-model Systems
Figure 1 and Figure 2 illustrates the Pareto-optimal lower envelopes for three configurations: single models, homogeneous multi-model ensembles composed of models from the same family, and heterogeneous multi-model ensembles composed of models from different families. Each point on the envelope corresponds to a Pareto-optimal configuration in terms of aggregated parameter count and oracle loss, while the solid curves represent nonlinear least-squares fits of the scaling law defined in Equation (6). Across all three settings, we observe a clear power-law relationship between loss and total parameter count. This result indicates that, despite increased structural complexity, multi-model systems exhibit regular and predictable scaling behavior analogous to classical single-model scaling laws.
Table 1: Parameter fitting results under different model collaboration settings. The loss floor $L_{\infty}$ is significantly reduced for model ensembles across families, signifying the benefit of multi-model collaboration.
Collaboration settings Raw points Pareto points $L_{\infty}$ AA $\alpha$
Single model 71 12 1765.0285 886.1545 0.3578
Model ensemble within a family 506 22 1780.0938 979.6996 0.8500
Model ensemble across families 1979 32 1625.5834 948.0984 0.5516
@@ 4.2 Quantitative Comparison of Scaling Exponents
To quantify differences in scaling efficiency, we fit the scaling law to the Pareto-optimal frontiers in each configuration. Table 1 reports the fitted parameters, including the scaling exponent $\alpha$ and the asymptotic loss floor $L_{\infty}$. The Single-model scaling yields an exponent of $\alpha=0.3578$, consistent with previously reported scaling behavior in language modeling. In contrast, multi-model systems achieve substantially larger scaling exponents. Heterogeneous ensembles composed of models from different families attain an exponent of $\alpha=0.5516$, indicating a significantly faster rate of loss reduction per unit increase in total parameters. Homogeneous ensembles within the same model family exhibit an even steeper exponent of $\alpha=0.8500$, reflecting rapid initial gains at smaller aggregated parameter budgets. These results demonstrate that multi-model collaboration enables more efficient utilization of parameters than single-model scaling, as evidenced by the consistently larger scaling exponents. Specifically, the fitted scaling laws corresponding to the three collaboration settings are given by:
$\mathcal{L}_{\text{single}}(P)$ $=\frac{886.1545}{P^{0.3578}}+1765.0285,$ (8)
$\mathcal{L}_{\text{same}}(P)$ $=\frac{979.6996}{P^{0.8500}}+1780.0938,$ (9)
$\mathcal{L}_{\text{diff}}(P)$ $=\frac{948.0984}{P^{0.5516}}+1625.5834,$ (10)
where $P$ denotes the total parameter count, and the subscripts indicate single-model, same-family ensemble, and cross-family ensemble settings, respectively.
@@ 4.3 Asymptotic Loss Floors
Beyond scaling speed, the fitted asymptotic loss floors $L_{\infty}$ reveal important qualitative differences between the three regimes. Single-model scaling exhibits an estimated loss floor of $L_{\infty}=1765.03$. Homogeneous multi-model ensembles yield a comparable loss floor of $L_{\infty}=1780.09$, suggesting limited asymptotic improvement when combining models within the same family. In contrast, heterogeneous ensembles achieve a substantially lower asymptotic loss floor of $L_{\infty}=1625.58$. This pronounced reduction indicates that cross-family collaboration enables access to complementary modeling capabilities that are unattainable by scaling a single architecture or combining closely related models.
4.4 Interpretation of Homogeneous versus Heterogeneous Scaling
Although homogeneous ensembles exhibit a steeper scaling exponent, their advantage is primarily confined to smaller parameter regimes. As the aggregated parameter budget increases, their performance rapidly saturates at a loss floor comparable to that of single models. This behavior suggests that homogeneous ensembles primarily reduce redundancy within a shared inductive bias, but do not fundamentally extend modeling capacity. Heterogeneous ensembles, by contrast, achieve both a faster scaling rate than single models and a significantly lower asymptotic loss floor. This combination indicates that architectural and training diversity enables persistent gains across scales, effectively shifting the performance frontier downward rather than merely accelerating convergence toward an existing limit.

### v2

@@ Abstract
Recent advances in large language models (LLMs) have been largely driven by scaling laws for individual models, which predict performance improvements as model parameters and data volume increase. However, the capabilities of any single LLM are inherently bounded. One solution originates from intricate interactions among multiple LLMs, rendering their collective performance surpasses that of any constituent model. Despite the rapid proliferation of multi-model integration techniques such as model routing and post-hoc ensembling, a unifying theoretical framework of performance scaling for multi-model collaboration remains absent. In this work, we propose the Law of Multi-model Collaboration, a scaling law that predicts the performance limits of LLM ensembles based on their aggregated parameter budget. To quantify the intrinsic upper bound of multi-model collaboration, we adopt a method-agnostic formulation and assume an idealized integration oracle where the total cross-entropy loss of each sample is determined by the minimum loss of any model in the model pool. Experimental results reveal that multi-model systems follow a power-law scaling with respect to the total parameter count, exhibiting a more significant improvement trend and a lower theoretical loss floor compared to single model scaling. Moreover, ensembles of heterogeneous model families achieve better performance scaling than those formed within a single model family, indicating that model diversity is a primary driver of collaboration gains. These findings suggest that model collaboration represents a critical axis for extending the intelligence frontier of LLMs.
@@ 1 Introduction
Motivated by these constraints, various multi-model integration approaches have been proposed (Chen et al., 2025b)—including dynamic routing (Qian et al., 2025), ensemble voting (Zhao et al., 2025), tool orchestration (Zhang et al., 2025), and agent-based collaboration (Tran et al., 2025; Chen et al., 2025a; Li et al., 2024; Fan et al., 2025)—which demonstrate superior performance and efficiency over individual models across diverse application scenarios. But these successes remain empirical and method-specific, justified through heuristics or task evaluations rather than a comprehensive understanding of why multi-model collaboration works. The result is a serious mismatch between theory and practice: although the scaling law explains the growth behavior of individual models, there is no similar framework for multi-model collaboration. Without theoretical grounding, we cannot compare multi-model systems to large single models rigorously or reason about their ultimate limits. This gap motivates a general law for multi-model collaboration, parallel to existing single-model scaling laws (Kaplan et al., 2020; Hoffmann et al., 2022; Kim et al., 2025).
@@
We propose the Law of Multi-model Collaboration, which characterizes the performance limit of multi-model collaborative systems as a function of the total number of parameters of all models involved. Rather than focusing on specific ensemble mechanisms, we adopt a method-agnostic view, treating multi-model systems as collective entities. This study operates within the AI Flow framework (An et al., 2025; Shao and Li, 2025), which advocates distributed intelligence integrating artificial intelligence with communication systems through collaborative interactions across heterogeneous models deployed at device, edge, and cloud levels (Yuan et al., 2025). Understanding fundamental performance limits of multi-model collaboration becomes a prerequisite for principled system design in this paradigm. Our formulation centers on a model pool of LLMs with total parameter count summed across the pool. For a given input distribution, we characterize the minimal achievable loss through model collaboration, defining a theoretical lower bound analogous to single-model scaling laws. To isolate intrinsic collaboration limits, we consider an idealized oracle integration where per-token cross-entropy loss equals the minimum achieved by any pool model. While unrealizable in practice, this oracle parallels optimality assumptions in prior scaling studies that abstract away algorithmic inefficiencies to reveal fundamental trends. We construct empirical performance frontiers by enumerating model combinations from a fixed pool, evaluating aggregated parameter budgets against achieved loss, and extracting Pareto-optimal points representing the theoretical performance envelope.
@@
Our analysis yields two key findings. First, multi-model systems exhibit clear power-law scaling between loss and total parameter count, following a similar scaling exponent to single models ($\alpha\approx 0.35$), but achieving a dramatically lower asymptotic loss floor—a 43% reduction from 2.21 to 1.25. This indicates that multi-model collaboration fundamentally shifts the achievable performance frontier downward, enabling access to performance regimes unattainable by any individual model regardless of scale. Second, scaling efficiency depends strongly on model diversity: pairwise analysis reveals that heterogeneous ensembles from different families (Yang et al., 2025; Grattafiori et al., 2024; Team et al., 2025) consistently outperform homogeneous same-family ensembles under equivalent parameter budgets, with cross-family combinations achieving substantially lower asymptotic loss floors. These findings imply model diversity constitutes a fundamental scaling dimension complementary to parameters, data, and compute. The Law of Multi-model Collaboration thus reframes ensemble methods from ad-hoc techniques into a systematic, theoretically grounded pathway for extending artificial intelligence frontiers. This work makes four contributions:
@@ 2.2 Text-level Loss under Heterogeneous Tokenization
To account for tokenization differences while maintaining comparability, we normalize the expected loss by the average token sequence length across all models and all test texts. Let $\bar{n}$ denote the average token sequence length over all models in the model pool and all texts in the evaluation dataset:
$\bar{n}=\frac{1}{|\mathcal{M}|\cdot|\mathcal{D}|}\sum_{M_{i}\in\mathcal{M}}\sum_{x\in\mathcal{D}}n_{i}(x).$ (2)
The expected loss of model $M_{i}$ over the distribution $\mathcal{D}$ is then given by
$\mathcal{L}(M_{i})\;=\;\frac{1}{\bar{n}}\mathbb{E}_{x\sim\mathcal{D}}\left[\mathcal{L}(M_{i},x)\right].$ (3)
This normalization enables fair comparison across models with heterogeneous tokenizers while preserving the sensitivity to modeling efficiency at the text level.
@@ 2.3 Oracle Multi-model Collaboration
That is, for each individual text, the oracle selects the single model within the pool that yields the lowest text-level loss (sum of token-level losses) under its own tokenizer. The expected oracle loss over the distribution $\mathcal{D}$ is then normalized by the average token sequence length $\bar{n}$:
$\mathcal{L}_{\mathrm{oracle}}(\mathcal{S})\;=\;\frac{1}{\bar{n}}\mathbb{E}_{x\sim\mathcal{D}}\left[\mathcal{L}_{\mathrm{oracle}}(\mathcal{S},x)\right].$ (5)
@@ 2.6 Law of Multi-model Collaboration
This law generalizes classical single-model scaling laws to the multi-model setting. Crucially, both the exponent $\alpha$ and the loss floor $\mathcal{L}_{\infty}$ depend on the composition of the model pool, with the trade-off between scaling rate and asymptotic performance being fundamentally shaped by model diversity. While homogeneous pools may exhibit steeper scaling exponents at smaller budgets, heterogeneous pools achieve substantially lower asymptotic loss floors, reflecting complementary modeling capabilities across architectural families.
@@ 3.2 Evaluation Dataset and Loss Measurement
All models are evaluated on a fixed held-out text dataset sampled from the distribution $\mathcal{D}$. For each model $M_{i}\in\mathcal{M}$ and each input text $x$, we compute the text-level loss $\mathcal{L}(M_{i},x)$ as defined in Equation (1), namely the sum of token-level cross-entropy losses under the model’s own tokenizer. To obtain the expected loss $\mathcal{L}(M_{i})$, we first compute the average of the summed token losses over all texts in the evaluation dataset, then normalize by the average token sequence length $\bar{n}$ across all models and all test texts. This normalization procedure yields a single-model performance baseline that is directly comparable across models despite heterogeneous tokenization schemes, while accounting for differences in tokenizer granularity.
@@ 3.3 Enumeration of Multi-model Combinations
To explore the space of multi-model collaboration while addressing the combinatorial explosion of possible subsets, we adopt a Pareto-pruning strategy. Directly enumerating all $2^{71}-1$ non-empty subsets of the model pool is computationally prohibitive. Instead, we employ an iterative procedure that constructs ensembles of size $k$ by incrementally adding models to Pareto-optimal ensembles of size k1k-1.
Specifically, we proceed as follows:
@@
    For $k=1$, evaluate all single models and extract the Pareto-optimal frontier based on parameter count and oracle loss.
@@
    For $k\geq 2$, consider only ensembles formed by adding one model from $\mathcal{M}$ to a Pareto-optimal ensemble of size k1k-1.
  3. 3.
    For each candidate ensemble $\mathcal{S}$, compute the aggregated parameter budget $P_{\mathrm{total}}(\mathcal{S})$ and the oracle loss $\mathcal{L}_{\mathrm{oracle}}(\mathcal{S})$ as defined in Equation (4).
  4. 4.
    Extract the Pareto-optimal frontier among all ensembles of size $k$, and proceed to $k+1$.
This iterative Pareto-pruning procedure enables us to explore ensembles ranging from 2 to 71 models while maintaining computational feasibility. The resulting configurations represent a representative sample of the multi-model scaling frontier, capturing the fundamental scaling trends without exhaustive enumeration.
To further investigate the role of model diversity, we also analyze pairwise combinations ($k=2$) separately, categorizing them into:
  1. 1.
    Homogeneous pairs, where both models belong to the same model family;
  2. 2.
    Heterogeneous pairs, where the models originate from different model families.
This partition enables a controlled comparison of scaling efficiency with and without architectural and training diversity at the pairwise level.
@@ 3.4 Pareto-optimal Frontier Extraction
Due to the large number of models and the wide variance in their capabilities, the raw set of evaluated points contains many dominated configurations. A configuration $\mathcal{S}$ is considered Pareto-optimal if no other configuration achieves both a lower oracle loss and a lower aggregated parameter budget.
To characterize the fundamental scaling behavior of multi-model collaboration, we extract Pareto-optimal frontiers for:
  • Single-model frontier: consisting of 12 Pareto-optimal models selected from the 71 base models;
  • Multi-model ensemble frontier: consisting of 3,146 Pareto-optimal ensembles selected from all explored combinations ranging from 2 to 71 models, obtained via the iterative Pareto-pruning procedure described above.
Additionally, to investigate the impact of model diversity on scaling efficiency, we separately extract Pareto-optimal frontiers from pairwise combinations:
  • Homogeneous pairs: 22 Pareto-optimal pairs selected from 506 same-family combinations;
  • Heterogeneous pairs: 32 Pareto-optimal pairs selected from 1,979 cross-family combinations.
The resulting Pareto frontiers represent the empirical lower envelopes of achievable loss under the oracle collaboration assumption.
Figure 1: Comparison between scaling performance under two settings: single model and multi-model ensemble. Solid curves denote nonlinear least-squares fits of the scaling formula, while points correspond to Pareto-optimal lower envelopes. Model ensembling requires fewer parameters to achieve a target loss and reduces the lower bound of the test loss significantly.
@@ 4.1 Scaling Behavior of Single and Multi-model Systems
Figure 1 compares the Pareto-optimal lower envelopes for single models and multi-model ensembles. The single-model frontier consists of 12 Pareto-optimal configurations selected from the 71 base models, while the multi-model ensemble frontier comprises 3,146 Pareto-optimal configurations obtained through the iterative Pareto-pruning procedure over ensemble sizes ranging from 2 to 71 models. Each point on the envelope corresponds to a Pareto-optimal configuration in terms of aggregated parameter count and oracle loss, while the solid curves represent nonlinear least-squares fits of the scaling law defined in Equation (7). Both settings exhibit a clear power-law relationship between loss and total parameter count, demonstrating that multi-model systems exhibit regular and predictable scaling behavior analogous to classical single-model scaling laws, despite their increased structural complexity.
Table 1: Fitted scaling law parameters for single-model and multi-model ensemble frontiers. The loss floor $L_{\infty}$ is significantly reduced for model ensembles, signifying the benefit of multi-model collaboration. The multi-model ensemble frontier aggregates all explored combinations from 2 to 71 models.
Setting Raw Points Pareto Points $L_{\infty}$ AA $\alpha$
Single model 71 12 2.21 1.11 0.3578
Model ensemble 3,146 3,146 1.25 2.02 0.3502
@@ 4.2 Quantitative Comparison of Scaling Exponents
To quantify differences in scaling efficiency, we fit the scaling law to the Pareto-optimal frontiers for both single-model and multi-model ensemble settings. Table 1 reports the fitted parameters, including the scaling exponent $\alpha$ and the asymptotic loss floor $L_{\infty}$. Single-model scaling yields an exponent of $\alpha=0.3578$, consistent with previously reported scaling behavior in language modeling. The multi-model ensemble frontier, which aggregates all explored combinations from 2 to 71 models, exhibits a comparable scaling exponent of $\alpha=0.3502$. While the scaling rates are similar, the critical distinction lies in the asymptotic behavior. The multi-model ensemble achieves a substantially lower loss floor of $L_{\infty}=1.25$, compared to $L_{\infty}=2.21$ for single models—a reduction of approximately 43%. This pronounced decrease indicates that multi-model collaboration fundamentally shifts the achievable performance frontier downward, enabling access to regions of the loss landscape unattainable by any individual model.
Specifically, the fitted scaling laws are given by:
$\mathcal{L}_{\text{single}}(P)$ $=\frac{1.11}{P^{0.3578}}+2.21,$ (9)
$\mathcal{L}_{\text{ensemble}}(P)$ $=\frac{2.02}{P^{0.3502}}+1.25,$ (10)
where $P$ denotes the total parameter count (in billions).
@@ 4.3 Asymptotic Loss Floors
The fitted asymptotic loss floors $L_{\infty}$ reveal the fundamental difference between single-model and multi-model scaling regimes. Single-model scaling exhibits an estimated loss floor of $L_{\infty}=2.21$, representing the theoretical lower bound achievable by any individual model architecture under the oracle assumption. In contrast, the multi-model ensemble frontier achieves a substantially lower asymptotic loss floor of $L_{\infty}=1.25$, representing a 43% reduction. This pronounced reduction indicates that multi-model collaboration enables access to complementary modeling capabilities that are unattainable by scaling a single architecture alone. The lower loss floor suggests that different models encode distinct hypotheses about the data distribution, and that their oracle combination effectively expands the representational capacity of the system. As the aggregated parameter budget increases, the multi-model ensemble asymptotically approaches a performance regime fundamentally inaccessible to any single model, regardless of its size.
4.4 The Role of Model Diversity: Homogeneous versus Heterogeneous Pairs
To investigate the impact of model diversity on collaboration efficiency, we separately analyze pairwise combinations, as illustrated in Figure 2. Among the 506 same-family pairs, 22 are Pareto-optimal; among the 1,979 cross-family pairs, 32 are Pareto-optimal.
The Pareto frontiers reveal distinct scaling behaviors, as summarized in Table 2. Homogeneous pairs (same-family combinations) exhibit a steep scaling exponent ($\alpha=0.8501$), reflecting rapid initial improvements at small parameter budgets. However, their performance quickly saturates at an asymptotic loss floor of $L_{\infty}=2.2314$, which is comparable to that of single models ($L_{\infty}=2.21$). This behavior suggests that combining architecturally similar models primarily reduces variance or redundancy within a shared inductive bias, rather than fundamentally extending representational capacity.
In contrast, heterogeneous pairs (cross-family combinations) exhibit a more moderate scaling exponent ($\alpha=0.5516$) but achieve a substantially lower asymptotic loss floor of $L_{\infty}=2.0377$. This indicates that architectural and training diversity is a primary driver of collaboration gains. While same-family pairs achieve rapid short-term improvements through redundancy reduction, cross-family pairs unlock complementary modeling capabilities that lower the fundamental performance limit. Models from different families encode distinct hypotheses about the data distribution, and their combination enables access to a broader region of the hypothesis space.
These findings demonstrate that model diversity is not merely a practical consideration but a fundamental factor governing the efficiency of multi-model scaling. The trade-off between steeper short-term scaling (homogeneous) and lower asymptotic limits (heterogeneous) highlights the importance of diversity in achieving sustained performance gains.
Table 2: Fitted scaling law parameters for pairwise model combinations. Same-family pairs exhibit steeper scaling exponents but saturate at loss floors comparable to single models, while cross-family pairs achieve substantially lower asymptotic loss through architectural diversity.
Combination Type Raw Pairs Pareto Pairs $L_{\infty}$ AA $\alpha$
Same-family pairs 506 22 2.2314 1.2281 0.8501
Cross-family pairs 1,979 32 2.0377 1.1885 0.5516